OpenStudy (anonymous):

Find a particular solution to the differential equation: y'' + 0y' - 1y = -1t^3

OpenStudy (anonymous):

ok, for this one, I would use the method of undetermined coefficients

OpenStudy (anonymous):

so it would take the form of A + Bt + Ct^2 + Dt^3

myininaya (myininaya):

y=A+Bt+Ct^2+Dt^3 y'=B+2Ct+3Dt^2 y''=2C+6Dt so we have 2C+6Dt-(A+Bt+Ct^2+Dt^3)=-1t^3

myininaya (myininaya):

2C-A=0 6D-B=0 -C=0 -D=-1

myininaya (myininaya):

If C=0, then 2(0)-A=0 implies A=0. If D=1, then 6(1)-B=0 implies -B=-6 implies B=6. So y=6t+t^3

myininaya (myininaya):

gotta make dinner peace